OpenStudy (jessie13):

Will Medal Answer ASAP!!!!! 1: What is the GCF of the numerator and denominator in the following fraction? 21m^2n/28mn^2 7m^2n^2 7mn 21m^3n^3 mn

OpenStudy (anonymous):

so whats the question?

OpenStudy (jessie13):

im putting it up right now!

OpenStudy (anonymous):

alrighty then.

OpenStudy (jessie13):

@paki

OpenStudy (jessie13):

@geny55 @Photon336

OpenStudy (photon336):

So this is what we have right @jessie13 \[\frac{ 21m ^{2}n }{ 28mn ^{2} }\]

OpenStudy (photon336):

so greatest common factor, is the largest number that can go into both the numerator and denominator

OpenStudy (jessie13):

1 dont get it!

OpenStudy (photon336):

Let's break this up into two parts

OpenStudy (photon336):

\[\frac{ 21 }{ 28 } * (\frac{ m ^{2}n }{ mn ^{2} })\]

OpenStudy (photon336):

do you know how to subtract powers?

OpenStudy (photon336):

whenever we divide exponents we subtract the exponents. so we will do that first \[\frac{ m ^{2}n}{ mn ^{2} } = Numerator[ m ^{2}-m ^{1} = m] denominator [n ^{1}-n ^{2}] = n ^{-1}\] we simplify the variables first. \[\frac{ m }{ n }\]

OpenStudy (photon336):

@jessie13 so now this is what we have we need to find the greatest common factor so it has to be the largest number that goes into both 21 and 28 \[\frac{ 21 }{ 28 } *(\frac{ m }{ n })\]

OpenStudy (jessie13):

wait the gcf is 7

OpenStudy (photon336):

@jessie13 yep :) here's why. and that's why I told you to do it that way yeah so